# ASTRODEEP-JWST: NIRCam-HST multiband photometry and redshifts for half a million sources in six extragalactic deep fields

### v1

@@ 4 Validation of photometry
To validate the accuracy of our catalogs, we compared our “optimal” catalogs (see Sect. 3.4) to other available ones, either published or obtained via private communication with the proprietary teams. We considered both archival catalogs based on HST observations (CANDELS and Astrodeep releases), and recent JWST catalogs from various research groups: P23 and Weaver et al. (2024) for ABELL2744, Finkelstein et al. (2023a) for CEERS, S. Finkelstein’s priv. comm. catalog for NGDEEP, and the JADES public catalogs for the two JADES fields, v2.0 for GOODS-South and v1.0 for GOODS-North (see Rieke et al., 2023) [^7] https://archive.stsci.edu/hlsp/jades.
@@
We cross-matched the catalogs using the equatorial coordinates of the detected sources with a searching radius of 0.35”. For the comparisons we only considered sources that have SNR $>5$ in both catalogs, and flag $<200$ in this work (i.e., well-behaved galaxies, non-blended in the detection stack; see Sect. 3.5).
@@
Since many details about the detection procedure vary significantly because of the different techniques adopted in the various works, we choose to focus on colors rather than on total fluxes, as they convey a more robust diagnostic on the accuracy of the estimated SED of the galaxies. The results are shown in Figs. 11 to 16 (the last panel for each field shows the relative difference in the total flux in F [^4] W for JWST catalogs and in F160W for HST catalogs, for the sake of completeness).
@@
The agreement is generally good, in particular with JWST catalogs and especially in the NIRCam bands. The larger differences are found with respect to the P23 catalog for ABEL2744 in the HST bluer bands, most likely because of the different PSF models and the local background subtraction introduced in this work. We also notice a systematic trend with respect to the JADES catalogs in the HST bands, with our colors typically becoming larger towards the faint end of the distribution; given the good agreement of the F [^4] W flux band estimates, this must be due to our fluxes being fainter in the HST bands. Concerning the comparisons with the archival HST catalogs, we notice an evident declining trend at the faint end of most plots, which must be due to the deeper sensitivity of the JWST data. Faint objects are now measured with more accuracy, whereas in the old catalog they often happened to have spurious positive fluxes (higher than their nominal errors, thus not being classified as upper limits) as a result of local noise fluctuations, especially in the less resolved redder bands ($Ks$ and IRAC), resulting in larger colors with respect to those measured in the high quality NIRCam images.
Furthermore, in the cases of ABELL2744 vs. the Astrodeep catalog by Merlin et al. (2016a) and NGDEEP vs. the CANDELS by Guo et al. (2013), the matched sources are very few, so the comparison is less significant.
@@
We feel confident that most of the discrepancies can be explained considering the differences in the adopted processing techniques, in particular concerning the apertures used to measure the fluxes, the PSF models, the background subtraction algorithms, the correction for galactic extinction. Similar discrepancies have been found among multi-wavelength catalogs in past efforts (see e.g. Stefanon et al., 2017); a more detailed analysis is beyond the scope of the present work.
@@ 5 Photometric redshifts
We estimated photometric redshifts on our “optimal” catalogs using the software packages zphot (Fontana et al., 2000) and EAzY (Brammer et al., 2008).
@@
For zphot we adopted Bruzual & Charlot (2003) templates, assumed exponentially declining Star Formation Histories (SFH) with timescale $\tau$ ranging from 0.1 to 15 Gyr, and included nebular emission lines according to Castellano et al. (2014) and Schaerer & de Barros (2009); metallicity can assume values of 0.02, 0.2, 1 and 2.5 times Solar and the age is allowed to vary from 10 Myr to the age of the Universe at a given redshift; finally, we adopted a Calzetti et al. (2000) extinction law with E(B-V) in the range 0–1.1.
@@
For EAzY we made three runs, one with the pre-defined set of templates eazy_v1.3 and the other two with the ones presented by Larson et al. (2023b) (we used the FSPS + Set 1 + Set 3 or “LyaReduced” and FSPS Set 1 + Set 4 or “Lya” combinations, as suggested by the authors; see https://ceers.github.io/LarsonSEDTemplates). So we end up with four redshift estimates, which we list in our final catalogs.
@@
As a final test on our photometry, we checked the accuracy of the photometric redshifts estimates by considering the subsample of sources having spectroscopic information from the literature. We matched these spectroscopic targets with our catalogs adopting a conservative searching radius of 0.3”. We considered the following spectroscopic samples: (i) NIRSpec: for ABELL2744, data from the programs GLASS-JWST-ERS-1324 (Treu et al., 2022), UNCOVER-GO-2561 (Bezanson et al., 2022) and GO-3073 (Castellano et al., 2024), using the Mascia et al. (2024) and Price et al. (2024, UNCOVER DR4) data releases, plus additional ones from our own on-going data analysis (Napolitano et al., in preparation); for CEERS, data from the program CEERS-ERS-1345 (Finkelstein et al., 2023a), retrieving them from the public DAWN JWST Archive [^8] https://dawn-cph.github.io/dja (see details in Heintz et al., 2024); for the JADES fields, data from the program JADES-GTO-1180 (Eisenstein et al., 2023), using the DR3 spectroscopic catalogs (D’Eugenio et al., 2024) released by the team [^9] https://jades-survey.github.io/scientists/data.html. Additionally, we also included the collection of spectra from various programs analysed and listed in Roberts-Borsani et al. (2024), and more data from the DAWN JWST Archive as available at the moment of writing this paper (July 2024); (ii) CANDELSz7 (Pentericci et al., 2018b) data, for JADES-GS, NGDEEP, and PRIMER; (iii) VANDELS DR4 (Pentericci et al., 2018a; McLure et al., 2018; Garilli et al., 2021) data, for PRIMER-UDS, JADES-GS and NGDEEP; (iv) the collection of ground based spectroscopy obtained with different instruments by different projects (VIMOS, Braglia et al. 2009, and MUSE, Mahler et al. 2018; Richard et al. 2021; Bergamini et al. 2023a, b) at the VLT, AAOmega on the Anglo-Australian Telescope, Owers et al. 2011, and of the HST WFC3/IR grism through the HST GO program GLASS (Treu et al., 2015; Schmidt et al., 2014) for the ABELL27444 field; (v) the collection compiled by the CANDELS collaboration and described in Kodra et al. (2023) for CEERS, JADES-GN PRIMER-UDS, PRIMER-COSMOS; (vi) the updated compilation of Merlin et al. (2021) for JADES-GS and NGDEEP; (vii) additional redshifts from Cowie et al. (2004); Reddy et al. (2006); Trump et al. (2009); van der Wel et al. (2016); Inami et al. (2017); Damjanov et al. (2018); Straatman et al. (2018); Scodeggio et al. (2018); Masters et al. (2019); Wisnioski et al. (2019); Urrutia et al. (2019); Ning et al. (2020); Jones et al. (2021); Pharo et al. (2022); Bacon et al. (2023) and the redshifts collected by Grazian et al. (2006); Wuyts et al. (2008); Xue et al. (2011); (viii) the MUSE (Schmidt et al., 2021; Rosani et al., 2020), zCOSMOS (Lilly et al., 2007), DEIMOS (Hasinger et al., 2018), VUDS (Tasca et al., 2017) catalogs for PRIMER-COSMOS. For our final comparisons we only included robust redshifts estimates on the basis of the various quality flags provided the authors, and took care to avoid repetitions. Whenever a spectroscopic target is listed in more than one survey and the inferred redshift estimates (flagged as robust) differ by more than $0.05\times(1+z_{avg})$, where $z_{avg}$ is the average spectroscopic redshift, we removed the target from the sample; if the difference is lower than this value but larger than 0.1, we considered the redshift, but we conservatively removed it from the statistics adopted to evaluate the accuracy of our photometric redshifts (see below).
@@
Figure 7: Comparative accuracy of photometric redshift estimates as a function of the SNR of the sources. The plot shows, for a random sample of 10,000 sources from the PRIMER-COSMOS field, the SNR at the wavelengths corresponding to the photometric bands included in the catalog, color-coding each point with the standard deviation of three photo-$z$ estimates. Objects with SNR¡10 in the red bands (and/or SNR¡1 in the blue bands) tend to have discordant photo-$z$ estimates. See text for more details)
@@
Since the templates do not include an AGN component, to correctly evaluate the accuracy of the inferred photometric redshifts we removed known AGNs from the catalogues, regardless of their nuclear contribution to the galaxy SED. To this aim, we took advantage of the samples presented by Harikane et al. (2023), Goulding et al. (2023), Maiolino et al. (2023), Kocevski et al. (2023), Larson et al. (2023a), Roberts-Borsani et al. (2024), Greene et al. (2024), Barro et al. (2024), based on JWST data, of X-ray catalogs (Nandra et al., 2015; Xue et al., 2016; Luo et al., 2017; Kocevski et al., 2018) by flagging as AGN all sources with 2-10 keV luminosity larger than $10^{42}$ erg/s or identified as AGN by the authors, and we removed spectroscopic targets flagged as AGNs from the VANDELS, zCOSMOS, DEIMOS, VUDS and COSMOS2020 (Weaver et al., 2022) datasets.
@@
Finally, we removed from the lists problematic sources using the flags described in Sect. 3.5. In particular, we excluded spurious detections, stars, sources that are saturated or at the boundary of the images - that is, we only kept sources with flag $<$400; in addition, we removed sources missing JWST photometry in more than three bands (i.e. having the tens figure in the flag larger than 3). With this criteria, we are left with a total sample of 16,263 spectra.
@@
Fig. 17 shows, for the total of the sources and for each field separately, the comparisons between the spectroscopic redshifts and the median values of the photo-$z$ estimates. The latter were computed excluding the EAzY “LyA” run (to avoid over-weighting the results from the runs using the templates by Larson, which are very similar in the two cases). Tab. 8 reports the full statistics for the four runs made with zphot and EAzY. We define $dz=|z_{phot}-z_{spec}|/(1+z_{spec})$, and $f_{outliers}$ as the percentage of objects with $dz<0.15$; we then compute mean, median, standard deviation, and NMAD (defined as 1.48$\times$×median($|dz|$)) of non-outliers. The four runs yield comparable statistics. The overall accuracy of the median estimates is good (NMAD 0.031, standard deviation 0.041 considering all fields together), and better than that from any individual zphot or EAzY run. There is a non-negligible fraction of outliers (6.3%) comparable to, albeit larger than, the one reached in recent efforts on multi-wavelength catalogs with a larger number of bands (and more sohisticated approaches, see e.g. Merlin et al., 2021). Looking at Fig. 17 we also note that the accuracy varies from field to field, with ABELL2744 and NGDEEP having the largest outlier fractions, most likely because of the effect of the cluster on photometric measurements for the former and of the low number of available spectroscopic redshifts for the latter.
@@
We report that we also tested delayed exponentially declining SFHs with the zphot code, and verified that our choice of standard exponentially declining models performs better in terms of agreement with spectroscopic redshifts.
@@
We point out that the photometric redshift estimates are prone to substantial uncertainties for the sources at low SNR. Fig. 7 exemplifies this by showing, for a random sub-sample of 10,000 sources in the PRIMER-COSMOS catalog, the value of the standard deviation of the photo-$z$ estimates obtained with three runs [^10] We computed the median and standard deviation of the estimates using the runs made with zphot, EAzY with eazy_v1.3 templates, and EAzY with Larson FSPS + Set 1 + Set 4 templates; we did not include the run with Larson FSPS + Set 1 + Set 3 templates since it yielded very similar results to the other one., as a function of the detection SNR in all bands: clearly, the discrepancy grows with decreasing SNR. Since we detected in the reddest bands, sources close to the detection limit in the F356W+F [^4] W stack typically become even fainter in bluer images, thus making their characterization very hard, as different codes and/or templates prefer different solutions for poorly constrained photometry.
@@
Then, as discussed in Sect. 3.5, taking the zphot runs as reference (given its overall slightly better performance) we also visually inspected all of the sources with estimated redshift above 10 (which were initially 7224). By doing so, not only did we exclude a further population of spurious detections, but we identified some systematic cases in which the redshift estimate was wrong due to unfortunate lack of relevant data: for example, in PRIMER-UDS some sources have been observed only with the LW NIRCam filters, and the missing information at wavelengths blue-ward of F277W caused a wrong identification of the Lyman break and therefore of the redshift. We marked these and similar cases with a negative sign before the flag value in the photometric redshifts catalogs. After these checks, we were left with 3101 objects with photo-$z>10$. This number is certainly an over-estimation. Indeed, only 922 have standard deviation of the estimates of the three codes lower than 0.5, again highlighting the challenging task of robustly constraining the photometric redshift for such faint, distant sources. Furthermore, some red low redshift interlopers must be present (see Arrabal Haro et al., 2023b; Harikane et al., 2024), and we deliberately only removed clearly spurious detections, not the uncertain or suspect ones – so certainly there are still fake sources polluting the sample. Comparing e.g. with the estimated surface density of high-$z$ objects by Finkelstein et al. (2023b, their Fig. 8), we would expect $\sim$20 galaxies at $z>10$ with F277W$<$28.5 in CEERS; we find 65 in the zphot sample (a value close to the expected number after correcting for completeness), but just 14 considering a restricted sample of sources having the standard deviation of the estimates from the three codes lower than 0.5.

### v3

@@ 4 Photometry validation
To validate the accuracy of our catalogues, we compared our ‘optimal’ catalogues (see Sect. 3.4) to other available ones, either published or obtained via private communication with the proprietary teams. We considered both archival catalogues based on HST observations (CANDELS and Astrodeep releases), and recent JWST catalogues from various research groups: P23 and Weaver et al. (2024) for ABELL2744, Finkelstein et al. (2023) for CEERS, S. Finkelstein’s priv. comm. catalogue for NGDEEP, and the JADES public catalogues for the two JADES fields, v2.0 for GOODS-South and v1.0 for GOODS-North (see Rieke et al., 2023) [^12] https://archive.stsci.edu/hlsp/jades, which also include photometry from the JEMS programme (Williams et al., 2023).
Figure 9: Example of a photometry validation plot, comparing the catalogues for the ABELL2744 field from this work and from Weaver et al. (2024). Relative errors, $\Delta c$, are given, namely, the colours measured in this work minus those in the reference catalogues), versus the second band magnitude in this work catalogue (e.g. F [^4] W for the F356-F [^4] W colour). The number of cross-matched sources after excluding those with S/N $<5$ in any of the two catalogues or flag $\geq 200$ in this work is also given; the blue line is the median of the distribution. Similar plots for all fields are available in the appendix.
@@
We cross-matched the catalogues using the equatorial coordinates of the detected sources with a searching radius of 0.35”. For the comparison, we only considered sources that have S/N $>5$ in both catalogues and flagged $<200$ in this work (i.e. well-behaved galaxies, non-blended in the detection stack; see Sect. 3.5).
@@
Since many details about the detection procedure vary significantly because of the different techniques adopted in the various works, we chose to focus on colours rather than on total fluxes, as they convey a more robust diagnostic on the accuracy of the estimated SED of the galaxies. The results are shown in Figs. D.1 to D.6. The last panel for each field shows the relative difference in the total flux in F [^4] W for JWST catalogues and in F160W for HST catalogues, for the sake of completeness. An example for one field (ABELL2744) is given in Fig. 9.
@@
The agreement is generally good, in particular with JWST catalogues and especially in the NIRCam bands. The larger differences are found with respect to the P23 catalogue for ABEL2744 in the HST bluer bands, most likely because of the different PSF models and the local background subtraction introduced in this work. We also noticed a systematic trend with respect to the JADES catalogues in the HST bands, with our colours typically becoming redder towards the faint end of the distribution; given the good agreement of the F [^4] W flux band estimates, this must be due to our fluxes being fainter in the HST bands. Concerning the comparisons with the archival HST catalogues, we notice an evident declining trend at the faint end of most plots, which must be due to the deeper sensitivity of the JWST data. Faint objects are now measured with more accuracy, whereas in the old catalogue, they often happened to have spurious positive fluxes (higher than their nominal errors, thus they were not classified as upper limits) as a result of local noise fluctuations; especially in the less resolved redder bands ($Ks$ and IRAC), resulting in larger colours with respect to those measured in the high quality NIRCam images. In the cases of ABELL2744 versus the Astrodeep catalogue by Merlin et al. (2016a) and NGDEEP versus the CANDELS by Guo et al. (2013), there are very few matched sources, so the comparison is less significant.
@@
We feel confident that most of the discrepancies can be explained considering the differences in the adopted processing techniques, particularly concerning the apertures used to measure the fluxes, the PSF models, the background subtraction algorithms, and the correction for galactic extinction. Similar discrepancies have been found among multi-wavelength catalogues in past efforts (see e.g. Stefanon et al., 2017). A more detailed analysis is beyond the scope of the present work.
@@ 5 Photometric redshifts
We estimated photometric redshifts on our ‘optimal’ catalogues using the software packages zphot (Fontana et al., 2000) and EAzY (Brammer et al., 2008). For zphot, we adopted templates from Bruzual & Charlot (2003) and we assumed exponentially declining star formation histories (SFHs), with timescales, $\tau$, ranging from 0.1 to 15 Gyr. We included nebular emission lines according to Castellano et al. (2014) and Schaerer & de Barros (2009). We considered metallicity values of 0.02, 0.2, 1, and 2.5 times Solar and the age was allowed to vary from 10 Myr to the age of the Universe at a given redshift. Finally, we adopted a Calzetti et al. (2000) extinction law with E(B-V) in the range of 0–1.1.
@@
For EAzY, we carried out three runs: one with the pre-defined set of templates eazy_v1.3 and the other two with the two sets from the ones presented by Larson et al. (2023b). We used the FSPS + Set 1 + Set 3 or ‘LyaReduced’ and FSPS Set 1 + Set 4 or ‘Lya’ combinations, as suggested by the authors [^13] https://ceers.github.io/LarsonSEDTemplates). So we end up with four redshift estimates, which we list in our final catalogues.
@@
As a final test of our photometry, we checked the accuracy of the photometric redshift estimates by considering the sub-sample of sources having spectroscopic information from the literature. We matched these spectroscopic targets with our catalogues adopting a conservative searching radius of 0.3”. We considered the following spectroscopic samples: (i) NIRSpec: for ABELL2744, data from the programmes GLASS-JWST-ERS-1324 (Treu et al., 2022), UNCOVER-GO-2561 (Bezanson et al., 2022) and GO-3073 (Castellano et al., 2024), using the Mascia et al. (2024) and Price et al. (2024, UNCOVER DR4) data releases, plus additional ones from our own on-going data analysis (Napolitano et al., 2024); for CEERS, data from the programme CEERS-ERS-1345 (Finkelstein et al., 2023), retrieving them from the public DAWN JWST Archive [^14] https://dawn-cph.github.io/dja (see details in Heintz et al., 2024); for the JADES fields, data from the programme JADES-GTO-1180 (Eisenstein et al., 2023), using the DR3 spectroscopic catalogues (D’Eugenio et al., 2024) released by the team [^15] https://jades-survey.github.io/scientists/data.html, and FRESCO-GO-1895 (Oesch et al., 2023; Meyer et al., 2024). Additionally, we also included the collection of spectra from various programmes analysed and listed in Roberts-Borsani et al. (2024), and more data from the DAWN JWST Archive as available at the moment of writing this paper (July 2024); (ii) CANDELSz7 (Pentericci et al., 2018b) data, for JADES-GS, NGDEEP, and PRIMER; (iii) VANDELS DR4 (Pentericci et al., 2018a; McLure et al., 2018; Garilli et al., 2021) data, for PRIMER-UDS, JADES-GS and NGDEEP; (iv) the collection of ground based spectroscopy obtained with different instruments by different projects (VIMOS, Braglia et al. 2009, and MUSE, Mahler et al. 2018; Richard et al. 2021; Bergamini et al. 2023a, b) at the VLT, AAOmega on the Anglo-Australian Telescope, Owers et al. 2011, and of the HST WFC3/IR grism through the HST GO programme GLASS (Treu et al., 2015; Schmidt et al., 2014) for the ABELL27444 field; (v) the collection compiled by the CANDELS collaboration and described in Kodra et al. (2023) for CEERS, JADES-GN PRIMER-UDS, PRIMER-COSMOS; (vi) the updated compilation of Merlin et al. (2021) for JADES-GS and NGDEEP; (vii) additional redshifts from Cowie et al. (2004); Reddy et al. (2006); Trump et al. (2009); van der Wel et al. (2016); Inami et al. (2017); Damjanov et al. (2018); Straatman et al. (2018); Scodeggio et al. (2018); Masters et al. (2019); Wisnioski et al. (2019); Urrutia et al. (2019); Ning et al. (2020); Jones et al. (2021); Pharo et al. (2022); Bacon et al. (2023) and the redshifts collected by Grazian et al. (2006); Wuyts et al. (2008); Xue et al. (2011); (viii) the MUSE (Schmidt et al., 2021; Rosani et al., 2020), zCOSMOS (Lilly et al., 2007), DEIMOS (Hasinger et al., 2018), and VUDS (Tasca et al., 2017) catalogues for PRIMER-COSMOS. For our final comparisons, we only included robust redshift estimates on the basis of the various quality flags provided by the authors and took care to avoid repetitions. Whenever a spectroscopic target was listed in more than one survey and the inferred redshift estimates (flagged as robust) differed by more than $0.05\times(1+z_{avg})$, where $z_{avg}$ is the average spectroscopic redshift, we removed the target from the sample; if the difference was lower than this value but larger than 0.1, we considered the redshift; however, we conservatively removed it from the statistics adopted to evaluate the accuracy of our photometric redshifts (see details below).
@@
Figure 10: Comparative accuracy of photometric redshift estimates as a function of the S/N of the sources. The plot shows, for a random sample of 20,000 sources from the PRIMER-COSMOS field, the S/N at the wavelengths corresponding to the photometric bands included in the catalogue, colour-coding each point with the standard deviation of the three photo-$z$ estimates used to compute the median photo-$z$ (see text for details). Objects with S/N¡10 in the red bands (and/or S/N¡1 in the blue bands) tend to have discordant photo-$z$ estimates. See text for more details.
@@
Since the templates do not include an AGN component, to correctly evaluate the accuracy of the inferred photometric redshifts, we removed known AGN from the catalogues, regardless of their nuclear contribution to the galaxy SED. To this aim, we took advantage of the lists based on JWST data presented by Harikane et al. (2023), Goulding et al. (2023), Maiolino et al. (2024), Kocevski et al. (2023), Larson et al. (2023a), Roberts-Borsani et al. (2024), Greene et al. (2024), Barro et al. (2024), and of X-ray catalogues (Nandra et al., 2015; Xue et al., 2016; Luo et al., 2017; Kocevski et al., 2018), by flagging as AGN all sources with 2-10 keV luminosity larger than $10^{42}$ erg/s or identified as AGN by the authors, and we removed spectroscopic targets flagged as AGN from the VANDELS, zCOSMOS, DEIMOS, VUDS, and COSMOS2020 (Weaver et al., 2022) datasets.
@@
Finally, we removed from the lists problematic sources using the flags described in Sect. 3.5. In particular, we excluded PLS and PSS, and sources that are saturated or at the boundary of the images; namely, we only kept sources with flags of $<$400. In addition, we removed sources lacking JWST photometry in more than three bands (i.e. having the second-to-last figure in the flag value larger than 3). Using these criteria, we were left with a total sample of 16,666 spectra.
@@
Figure 20 shows the global comparison between the spectroscopic redshifts and the median values of the photo-$z$ estimates. The latter were computed excluding the EAzY ‘LyA’ run, to avoid over-weighting the results from the runs using the templates by Larson, which are very similar in the two cases. Similar plots for each field individually are shown in Appendix E; Table E.1 reports the full statistics for the four runs made with zphot and EAzY. We define $dz=|z_{phot}-z_{spec}|/(1+z_{spec})$, and $f_{outliers}$ as the percentage of objects with $dz<0.15$. We then compute mean, median, standard deviation, and NMAD (defined as 1.48$\times$×median($|dz|$)) of non-outliers. The four runs yield comparable statistics. The overall accuracy of the median estimates is good (NMAD 0.031, standard deviation 0.041 considering all fields together) and better than that from any individual zphot or EAzY run. There is a non-negligible fraction of outliers (6.3%) comparable to (albeit larger than) the one reached in recent efforts on multi-wavelength catalogues with a larger number of bands (and more sohisticated approaches; see e.g. Merlin et al., 2021). Looking at Fig. E.1, we also note that the accuracy varies from field to field, with ABELL2744 and NGDEEP having the largest outlier fractions. This is most likely because of the effect of the cluster on photometric measurements for the former and of the low number of available spectroscopic redshifts for the latter.
[…]
@@
We report that we also tested delayed exponentially declining SFHs with the zphot code. We verified that our choice of standard exponentially declining models performs better in terms of agreement with spectroscopic redshifts.
@@
We point out that the photometric redshift estimates are prone to substantial uncertainties for the sources at low S/N. Figure 10 exemplifies this by showing the S/N as a function of the band wavelength for a random sub-sample of 20,000 sources in the PRIMER-COSMOS catalogue, colour-coding the points by the value of the standard deviation of the three photo-$z$ estimates used to compute the median value. Clearly, the discrepancy between the estimates grows with decreasing S/N. Since we detected in the reddest bands, sources close to the detection limit in the F356W+F [^4] W stack typically become even fainter in bluer images, thus making their characterisation very hard. This is mainly because different codes and/or templates prefer different solutions for poorly constrained photometry.
@@
Then, as discussed in Sect. 3.5, taking the zphot runs as reference (given its overall slightly better performance), we also visually inspected all of the sources with estimated redshift above 10 (which were initially 7123). By doing so, not only did we exclude a further population of spurious detections, but we identified some systematic cases where the redshift estimate was wrong due to unfortunate lack of relevant data. For example, in PRIMER-UDS some sources have been observed only with the LW NIRCam filters; thus, the missing information at wavelengths blue-ward of F277W caused an erroneous identification of the Lyman break and therefore of the redshift. We marked these and similar cases with a negative sign before the flag value in the photometric redshift catalogues. After these checks, excluding sources flagged as spurious we were left with 3068 objects with photo-$z>10$. This number is certainly an over-estimation. Indeed, only 798 have a standard deviation of the estimates of the three codes lower than 0.5; again highlighting the challenging task of robustly constraining the photometric redshift for such faint, distant sources. Furthermore, some red low redshift interlopers must be present (see Arrabal Haro et al., 2023b; Harikane et al., 2024) and we deliberately only removed clearly spurious detections, not the uncertain or suspect ones. Thus, there are certainly still a number of fake sources polluting the sample. Comparing, for instance, with the estimated surface density of high-$z$ objects by Finkelstein et al. (2024, as seen in their Fig. 8), we would expect $\sim$20 galaxies at $z>10$ with F277W$<$28.5 in CEERS; we found 59 in the zphot sample (a value close to the expected number after correcting for completeness), but just 9 considering a restricted sample of sources having a standard deviation of the estimates from the three codes that is lower than 0.5.